Sketch the region enclosed by the given curves. Decide whether to integrate with respect to x or y. Then find the area of the region.

2y = 4sqrt(x), y=5 and 2x + 2y =6.



1
Expert's answer
2022-04-29T10:10:35-0400
"2x+2y=6=>x=-y+3"

"2y = 4\sqrt{x}=>x=\dfrac{1}{4}y^2, y\ge0"

"-y+3=\dfrac{1}{4}y^2, y\ge0"

"y^2+4y-12=0, y\ge0"

"(y-2)(y+6)=0, y\ge0"

Then "y=2."


"Area=A=\displaystyle\int_{2}^{5}(\dfrac{1}{4}y^2-(-y+3))dy"

"=\bigg[\dfrac{1}{12}y^3+\dfrac{1}{2}y^2-3y\bigg]\begin{matrix}
 5 \\
 2
\end{matrix}"

"=\dfrac{125}{12}+\dfrac{25}{2}-15-\dfrac{8}{12}-2+6"

"=\dfrac{45}{4}({units}^2)=11.25({units}^2)"

"Area=11.25" square units.